# A hidden Markov model for statistical arbitrage in international crude oil futures markets

### v2

@@ 4.2. Testing the strategies
We now analyze over the test sample (from $t_{B}=$ 07/01/2022 until $T=$ 06/30/2023) the five different statistical arbitrage strategies described in Section 3.1: the plain Vanilla (PV), the probability interval (ProbI), the prediction interval (PredI), the realized increment (RI), the predicted increment (PI), along with the buy-and-hold S&P500 and the crude oil ETF strategies. For each strategy we compute the opening/closing signals, the annualized return according to (3.2) and the annualized Sharpe ratio according to (3.3).
@@
Figure 6 illustrates how the signals differ across the five statistical arbitrage strategies.
@@
Figure 6. Opening/closing signals of the statistical arbitrage strategies described in Section 3.1 over the test sample of Section 4.1.
@@
We immediately notice that the number of trades over the test sample varies significantly across the different strategies. Excluding the PV strategy that entails always an open position as soon as $S_{t}\neq 0$, we document a minimum of five trades with the ProbI strategy and a maximum of twenty-nine trades with the PredI strategy. Moreover, also the length of the open positions varies considerably: indeed, we generate positions that are opened on a given day and closed on the following day together with positions that remain open for several months.
PV ProbI PredI RI PI S&P ETF
$R$ 13.03% 56.37% 72.77% -1.43% 22.41% 16.43% -25.46%
$SR$ 0.6082 1.0849 1.0108 0.4440 0.6836 0.8868 -0.6299
Table 4. Annualized performances with $c=80$bps over the test sample of Section 4.1 including the Brent, the Shanghai and the WTI futures contracts.
Table 4 summarizes the return ($R$) and Sharpe ratio ($SR$), as described in Section 3.2, assuming $c=80$bps. As mentioned in Section 3.2, this relatively high level of transaction costs yields a conservative assessment of the performance of our statistical arbitrage strategies involving the Shanghai futures. First of all, we notice that the ETF on the crude oil futures market performs poorly on the test sample. This is coherent with the global decline of crude oil futures prices over the test sample which is clear from Figure 1. Moving to statistical arbitrage strategies, we find that only the PV and RI strategies deliver a return lower than the S&P strategy, which entails no transaction costs. On the contrary, the ProbI, PredI and PI strategies outperform the S&P strategy and deliver significant returns, overcoming the considerable amount of transaction costs assumed in our analysis. In particular, it turns out that the forward-looking strategies PredI and PI that are based on the stochastic model introduced in Section 2 outperform the simpler ProbI and RI strategies that rely only on past observations of the spread. We can therefore conclude that our OU-HMM model seems to have promising applications for optimal investment in international crude oil futures markets.
@@ 4.2.1. Risk analysis
Statistical arbitrage strategies might involve a considerable financial risk. In particular, when opening a position, the investor does not know when the futures prices will realign or whether convergence will ever be reached. As a consequence, the investor might find herself stuck in a costly position to maintain (due to margin calls) and that does not generate any profit for a prolonged period of time. In this section, we assess the riskiness of the statistical arbitrage strategies by means of a Monte Carlo analysis.
First of all, we compute the out-of-sample daily Value-at-Risk (VaR henceforth) of the strategies described in Section 3.1. Moreover, we exploit the simulations that are required for VaR estimation in order to derive the empirical distribution of the annualized returns of our strategies. This is done by implementing the following successive steps:
Figure 7. A simulated out-of-sample trajectory of the spread.
  1. (1)
    we estimate the cointegration relationship over the training sample;
  2. (2)
    we filter the resulting spread over the training sample following Section 4.1.3;
  3. (3)
    at the breaking date $t_{B}$ we retrieve the optimal estimated number of states $N$ of the Markov chain and the related point estimates of the parameters of our OU-HMM model;
  4. (4)
    for $i=1,...,NSim$:
    • we simulate over the test sample (i.e. from $t_{B}$ to $T$) a trajectory of the latent Markov chain X with the optimal $N$ found before;
    • we simulate over the test sample the spread process following (2.3) and the realizations of X simulated at the previous step using the parameters obtained at the end of the training sample. Figure 7 shows a simulated trajectory of the spread;
    • we compute the series of the returns according to (3.1) of the different statistical arbitrage strategies described in Section 3.1;
    • we compute the daily VaR at the confidence levels 1%, 5% and 10%;
    • we compute the overall performance of the strategies in terms of annualized return (3.2) and Sharpe ratio (3.3);
  5. (5)
    we average out the VaRs obtained along all the $NSim$ trajectories;
  6. (6)
    we compute the Kernel Density Estimator (KDE henceforth) of the annualized returns obtained along all the $NSim$ trajectories.
PV ProbI PredI RI PI S&P ETF
$VaR_{99\%}$10.21% -10.86% -10.16% -7.93% -7.62% -2.80% -5.29%
$VaR_{95\%}$5.88% -5.49% -5.80% -3.66% -3.69% -1.72% -3.80%
$VaR_{90\%}$4.04% -3.35% -3.87% -1.94% -2.12% -1.29% -3.13%
@@
Table 5. Monte Carlo estimates of the daily VaRs of daily returns over the test sample of Section 4.1 including the Brent, the Shanghai and the WTI futures contract ($NSim=1000$; the average t-statistics is equal to almost 6).
Table 5 displays the Monte Carlo estimates of the daily VaRs at the usual confidence levels. As expected, the attractive performances of statistical arbitrage opportunities involving the Shanghai futures reported in Table 4 are mitigated by a considerable level of risk, if compared to the simple S&P strategy. This implies that an arbitrageur willing to engage in statistical arbitrage strategies involving the Shanghai futures has to be ready to face possibly severe drops of the market value of her portfolio which may call for extra margins to be posted.
@@
Figure 8. Monte Carlo-based KDEs of annual returns over the test sample of Section 4.1 including the Brent, the Shanghai and the WTI futures contract ($NSim=1000$).
Figure 8 displays the Monte Carlo-based KDEs of the annual returns of the five strategies over the test sample. As we can see, all densities are positively skewed, with a sizeable amount of probability mass associated to positive returns. This provides evidence that statistical arbitrage strategies that are designed by exploiting the OU-HMM model introduced in Section 2 yield profitable investment opportunities.
@@ 4.3. Choice of the underlying futures contracts
In order to assess the relevance of the results reported in the previous subsection, we test the same statistical arbitrage strategies using different choices of futures contracts. First, we follow the traditional approach to pairs trading by using strategies that involve only two futures contracts instead of three. Then, we consider an alternative triplet of securities, replacing the recently introduced Shanghai futures with the Dubai crude oil futures, a well-established benchmark for crude oil.
The results of this analysis are reported in Table 6. For comparison purposes, the first row replicates the data from Table 4, while the performances of the benchmark passive strategies, S&P (R: 16.43%, SR: 0.8868) and ETF (R: -25.46%, SR: -0.6299), are omitted. Transaction costs are set at $c=80$bps when trading the Shanghai futures and reduced to $c=20$bps when only well established crude oil futures (Brent, WTI, and Dubai) are used. As discussed in Subsection 3.2, $c=80$bps is likely an overestimate. However, we prefer to maintain this conservative figure to account for any potential friction an investor might encounter when trading the recently introduced Shanghai futures. In contrast, $c=20$bps represents a lower bound of transaction costs, which is more plausible when trading well-established futures contracts.
As shown in Table 6, all the proposed statistical arbitrage strategies, except for the RI strategy, perform best when trading the Brent, Shanghai, and WTI futures, compared to any other combination involving these three or the Dubai futures, and despite the much higher transaction costs considered for the Shanghai futures. The second-best performing set of futures to trade is the simple Brent-WTI pair, which delivers positive returns for all strategies. While the triplet Brent-Shanghai-WTI is the most profitable, when replacing the recently introduced Shanghai futures with the Dubai one, the performance of our statistical arbitrage strategies deteriorate, despite the lower transaction costs. This might be explained by the fact that, as soon as the underlying assets become widespread, statistical arbitrage opportunities tend to disappear, as already noted in the literature (see, e.g., [Do and Faff, 2010]).
Finally, it is noteworthy that, over this specific training period, the WTI does not appear to cointegrate with neither the Shanghai nor the Dubai. Although unusual, this can happen occasionally, as can also be seen from Figure 10, and might indicate some persistent deviation of the WTI from the other crude oil futures.
$c$ PV ProbI PredI RI PI
@@
$R$ 80bps 13.03% 56.37% 72.77% -1.43% 22.41%
$SR$ 0.6082 1.0849 1.0108 0.444 0.6836
@@
$R$ 80bps -59.05% 36.29% -51.21% -24.11% -50.73%
$SR$ 0.003 0.8127 0.1319 -1.0759 0.1704
Shanghai - WTI
No statistical evidence of cointegration
Brent - WTI
$R$ 20bps 11.39% 10.39% 10.47% 4.88% 11.4%
$SR$ 0.8891 0.8314 0.8278 0.5404 0.8895
@@
$R$ 20bps 6.45% 7.21% 7.22% $0$60%
$SR$ 0.6195 0.7115 0.6852 0.0363 -0.0299
Brent - Dubai
$R$ 20bps 1.24% 5.65% -0.21% -0.61% -4.64%
$SR$ 0.171 0.5978 0.0295 -0.2278 -1.2266
WTI - Dubai
No statistical evidence of cointegration
Table 6. Annualized performances over the test sample including different combinations of the Brent, the Dubai, the Shanghai and the WTI futures contracts.

### v3

@@ 4.2. Testing the strategies
We now analyze over the test sample (from $t_{B}=$ 07/01/2022 until $T=$ 06/30/2023) the five statistical arbitrage strategies described in Section 3.1: the plain Vanilla (PV), the probability interval (ProbI), the prediction interval (PredI), the realized increment (RI), the predicted increment (PI). For all strategies except the PV benchmark, we consider several standard values of the bandwidth level $\alpha$. As a comparison, we also include the buy-and-hold S&P500 and the crude oil ETF strategies.
@@
Figure 6 illustrates how the trading signals differ across the five statistical arbitrage strategies, for the baseline value $\alpha=0.20$ of the bandwidth level.
@@
Figure 6. Opening/closing signals of the statistical arbitrage strategies described in Section 3.1, with bandwith level $\alpha=0.20$, over the test sample described in Section 4.1.
@@
We immediately notice that the number of trades over the test sample varies significantly across the different strategies. Excluding the PV strategy that entails always an open position as soon as $S_{t}\neq 0$, we document a minimum of twelve trades with the PredI strategy and a maximum of thirty-five trades with the PI strategy. Moreover, also the length of the open positions varies considerably: indeed, we observe trading positions that are opened on a given day and closed on the following day together with positions that remain open for several months.
PV ProbI PredI RI PI Ex. S&P Ex. ETF
$\alpha=0.20$
$R$ -0.0091 0.1043 0.1518 -0.0244 0.0230 0.1174 -0.2847
SRSR -0.0025 0.8335 1.1792 -0.1508 0.2359 0.6729 -0.7436
$p$-val. SR test 0.5125 0.0765 0.0615 0.635 0.3790 0.2360 0.8710
$N_{\mathrm{TW}}$ 37 24 12 30 35
mean $R^{\mathrm{TW}}$ -0.0002 0.0042∗∗ 0.0120∗∗ -0.0008 0.0007
$p$-val. $R^{\mathrm{TW}}$ 0.9053 0.0251 0.0467 0.6547 0.7013
% $R^{\mathrm{TW}}_{j}>0$ 0.5135 0.6667 0.8333 0.4000 0.5714
$p$-val. WRC 0.0650 best from WRC: PredI
ProbI PredI RI PI ProbI PredI RI PI
$\alpha=0.30$ $\alpha=0.25$
$R$ 0.1055 0.1394 -0.0471 0.0174 0.1043 0.1668 -0.0432 0.0243
SRSR 0.8151 1.1010 -0.3137 0.1955 0.8335 1.2562∗∗ -0.2909 0.2458
$p$-val. SR test 0.0795 0.0605 0.7270 0.3995 0.0765 0.0360 0.7170 0.3715
$N_{\mathrm{TW}}$ 26 18 35 36 24 15 33 35
mean $R^{\mathrm{TW}}$ 0.0039∗∗ 0.0073∗∗ -0.0013 0.0005 0.0042∗∗ 0.0104∗∗∗ -0.0013 0.0007
$p$-val. $R^{\mathrm{TW}}$ 0.0496 0.0202 0.4320 0.7639 0.0251 0.0075 0.4702 0.6852
% $R^{\mathrm{TW}}_{j}>0$ 0.6538 0.8333 0.4000 0.5556 0.6667 0.8667 0.3939 0.5714
$p$-val. WRC 0.0745 best from WRC: PredI 0.0360 best from WRC: PredI∗∗
$\alpha=0.15$ $\alpha=0.05$
$R$ 0.0572 0.1632 0.0206 0.0416 0.0559 0.0729 -0.0020 0.1994
SRSR 0.5171 1.3032∗∗ 0.2373 0.3685 0.7556 0.7244 -0.0319 1.3136∗∗
$p$-val. SR test 0.1520 0.0180 0.3670 0.2925 0.1030 0.1250 0.5290 0.0135
$N_{\mathrm{TW}}$ 18 10 24 34 7 5 11 27
mean $R^{\mathrm{TW}}$ 0.0031 0.0153∗∗ 0.0009 0.0013 0.0079 0.0142 -0.0001 0.0068∗∗
$p$-val. $R^{\mathrm{TW}}$ 0.2251 0.0110 0.6727 0.4944 0.1657 0.0551 0.9720 0.0136
% $R^{\mathrm{TW}}_{j}>0$ 0.6667 0.90000 0.5417 0.6176 0.7143 1.0000 0.5455 0.7407
$p$-val. WRC 0.0335 best from WRC: PredI∗∗ 0.0110 best from WRC: PI∗∗
Table 5. Performances of the statistical arbitrage strategies for different bandwidth levels $\alpha$ over the test sample of Section 4.1 considering the Brent, the Shanghai and the WTI futures. Statistical significance is denoted by , ∗∗ and ∗∗∗, corresponding to the 10%, 5% and 1% levels, respectively.
Table 5 summarizes the performance of all the strategies introduced in Section 3.1, considering different values of the bandwidth level $\alpha$ and the transaction cost parameters reported in Table 1. Note that the PV, Ex. S&P, and Ex. ETF strategies do not depend on $\alpha$: accordingly, their performance measures are reported only once in the table.
To assess the statistical significance of the performance of the proposed strategies, Table 5 reports, alongside the performance metrics considered in Section 3.2, the results of several statistical tests. For each statistical arbitrage strategy, the following quantities are reported:
  • the overall annualized return $R$ over the test sample, as defined in (3.2);
  • the annualized Sharpe Ratio $SR$, defined in (3.4);
  • the $p$-value of the test [^7] The test is implemented using the fixed-size circular block bootstrap considered in [Ledoit and Wolf, 2008], based on 2,000 bootstrap replications with 20 blocks each. for the statistical significance of the Sharpe Ratio, as proposed in [Ledoit and Wolf, 2008], conducted against the null hypothesis $H_{0}:SR=0$, which is appropriate since the performance of all strategies is evaluated in terms of excess returns;
  • the total number $N_{\mathrm{TW}}$ of completed trades over the entire test sample;
  • the sample mean of the trading-window returns $R^{\mathrm{TW}}=\{R^{\mathrm{TW}}_{1},\ldots,R^{\mathrm{TW}}_{N_{\mathrm{TW}}}\}$;
  • the $p$-value of the standard $t$-test for the null hypothesis that the mean of $R^{\mathrm{TW}}$ is zero;
  • the percentage of strictly positive trading-window returns in $R^{\mathrm{TW}}$;
  • the $p$-value of the White Reality Check (WRC) [^8] The test is implemented using the stationary bootstrap of [Politis and Romano, 1994], with 2,000 bootstrap replications and an average block length of 20. proposed by [White, 2000] and implemented following [Sullivan et al., 1999]. This test evaluates whether, within a set of strategies backtested on the same dataset, at least one strategy significantly outperforms both the others and a benchmark, here again taken to be a zero-return investment. Along with the WRC $p$-value, we also report the best-performing strategy in our sample.
First of all, we notice that the ETF on the crude oil futures market performs poorly on the test sample. This is coherent with the global decline of crude oil futures prices over the test sample which is clear from Figure 1. In contrast, the S&P performed relatively well, but its Sharpe Ratio is not statistically significant at any conventional confidence level.
Overall, the results reported in Table 5 indicate a clear advantage of the forward-looking strategies based on the proposed AR-HMM model, namely PredI and PI, over the model-free backward-looking strategies ProbI and RI. Across all considered values of the bandwidth level, PredI consistently yields the highest annualized returns and Sharpe ratios among the statistical arbitrage strategies, often attaining statistical significance according to the Sharpe Ratio test. The PI strategy also performs strongly, particularly for smaller values of $\alpha$, where it combines high returns with a limited number of trades and favorable trading-window statistics. In contrast, the backward-looking ProbI strategy exhibits positive but generally weaker performance, while the RI strategy fails to deliver consistent profitability and is never statistically significant. The superiority of the forward-looking approaches is further confirmed by the White Reality Check, which repeatedly identifies PredI (and, for the smallest value of $\alpha$, PI) as the best-performing strategy within the set, with rejection of the null hypothesis of no superior strategy at conventional significance levels. Taken together, these results suggest that explicitly modeling the joint dynamics of the spread and its latent regimes yields economically and statistically meaningful gains relative to purely distributional backward-looking alternatives.
Finally, we observe that the magnitude of these returns is consistent with the results in [Hain et al., 2018], that document that this kind of statistical arbitrage strategies involving commodity futures yield excess returns of about 6%-8% annually.
Remark 4.2.
It is worth noting that, for all statistical arbitrage strategies tested in our study, the series of daily returns include extended periods in which no position is open and returns are therefore equal zero. The resulting dilution of the unconditional daily mean may attenuate the Sharpe ratio relative to the profitability realized during active trading windows, potentially reducing the power of statistical Sharpe ratio tests based on daily returns. This issue further motivates the interest of performing a complementary analysis based on trading-window returns, as introduced in Section 3.2.
Remark 4.3 (Robustness of trading strategy performance to filtering initialization).
Table 6 provides evidence on the robustness of the performance of the PredI and PI strategies with respect to the initialization of the filtering algorithm. For conciseness, we report the annualized return $R$, the Sharpe ratio $SR$, and the mean trading-window return $R^{\mathrm{TW}}$. Across 100 runs of the filter-based EM algorithm with randomized initial conditions as described in Remark 4.1, both strategies exhibit limited dispersion in all reported performance measures. The standard deviations of $R$, SRSR and $R^{\mathrm{TW}}$ are small relative to their corresponding means, resulting in consistently large $t$-ratios. This pattern indicates that the observed profitability is not driven by specific initialization choices or by convergence to isolated local optima of the likelihood, but instead reflects stable features of the underlying spread modeling and filtering framework.
PredI PI
$R$ SRSR mean $^{\mathrm{TW}} RR$ SRSR mean $R^{\mathrm{TW}}$
baseline 0.1518 1.1792 0.012 0.023 0.2359 0.0007
mean 0.1679 1.2718 0.0125 0.0374 0.3406 0.0012
std. dev. 0.0349 0.254 0.0033 0.0094 0.0706 0.0005
$t$-ratio 4.808 5.008 3.8159 4.0001 4.8235 2.484
Table 6. Robustness of the performance of the PredI and PI strategies with respect to the initialization of the filtering algorithm. Reported means and standard deviations are computed from 100 runs of the algorithm with random initializations, as described in Remark 4.1.
@@ 4.2.1. Risk analysis
Statistical arbitrage strategies might involve a considerable financial risk. In particular, when opening a position, the investor does not know when the futures prices will realign or whether convergence will ever be reached. As a consequence, the investor might find herself stuck in a costly position to maintain (due to margin calls) and that does not generate profits for a prolonged period of time. In this section, we assess the riskiness of the statistical arbitrage strategies by means of a bootstrap-based analysis. We conduct the analysis using the baseline bandwidth level $\alpha=0.20$, an intermediate specification within the range considered above that balances trading frequency and transaction costs, as discussed in Section 3.1.
To provide a comprehensive assessment of the risk profile of the trading strategies described in Section 3.1, and to overcome the limitations of the Sharpe ratio, which depends solely on the first two moments of the return distribution, we adopt a resampling-based approach. Specifically, we employ the stationary bootstrap of [Politis and Romano, 1994] with 2,000 replications and an expected block length of 20 to account for serial dependence in daily returns. Using this procedure, we compute out-of-sample risk measures at daily frequency, including the left-tail Value-at-Risk (VaR), the corresponding Expected Shortfall (ES) and the Maximum Drawdown (MDD). In addition, the bootstrap resampling scheme allows us to recover the empirical distribution of annual returns $R$. This enables a full characterization of the distributional properties of strategy performance beyond location and scale measures.
PV ProbI PredI RI PI Ex. S&P Ex. ETF
VaR99% -2.5078% -2.5078% -2.5078% -2.6663% -2.5078% -2.8106% -5.3009%
VaR95% -1.5197% -1.3472% -1.3178% -1.4422% -1.5197% -1.7337% -3.8152%
VaR90% -1.1033% -1.0294% -0.8052% -1.0294% -1.1033% -1.3019% -3.1380%
ES99% -2.7297% -2.7297% -2.7297% -2.8093% -2.7297% -3.5073% -6.9106%
ES95% -1.8981% -1.8324% -1.8155% -1.896% -1.8981% -2.435% -5.0296%
ES90% -1.5848% -1.5259% -1.4400% -1.5728% -1.5848% -1.9961% -4.2791%
MDD -10.1668% -6.3788% -5.7282% -10.7311% -9.3580% -15.5447% -39.4747%
@@
Table 7. Bootstrapped estimates of the daily VaRs/ESs/MDD of daily returns over the test sample of Section 4.1 including the Brent, the Shanghai and the WTI futures contracts.
The results reported in Table 7 indicate that the proposed strategies exhibit substantially lower tail risk and drawdown relative to the excess return benchmarks, in particular when compared to the Ex. ETF strategy, which displays larger losses and significantly deeper drawdowns. By contrast, the proposed statistical arbitrage strategies tend to generate less volatile returns. This aligns with the trading signals shown in Figure 6, where, apart from one long position shortly after the start of the test period, most positions are held for relatively short durations, allowing for quick convergence and the realization of positive profits. While mild differences across strategies emerge at conventional levels, at the 99% level the scarcity of extreme observations limits the empirical discrimination of downside risk measures.
@@
Figure 7. Bootstrapped KDEs of annual returns over the test sample of Section 4.1 including the Brent, the Shanghai and the WTI futures contracts.
Figure 7 reports the bootstrap-based kernel density estimates of the annualized returns of the five strategies introduced in Section 3.1 over the test sample. The dispersion and location of the distributions differ across strategy classes. In particular, the plain vanilla (PV) strategy exhibits a comparatively symmetric distribution concentrated around its central tendency. We can observe that for the ProbI and PredI strategies the estimated densities are centered around positive returns. The density of the PredI strategy is moderately shifted to the right relative to that of its backward-looking counterpart, ProbI, suggesting a tendency toward higher returns. By contrast, the distributions of the RI and PI strategies display noticeable differences in shape and dispersion, with neither strategy dominating the other across the support of the distribution.
@@ 4.3. Choice of the underlying futures contracts
In order to assess the relevance of the results reported in the previous subsection, we test the same statistical arbitrage strategies using different choices of futures contracts. First, we follow the traditional approach to pairs trading by using strategies that involve only two futures contracts instead of three. Then, we consider an alternative triplet of securities, replacing the recently introduced Shanghai futures with the Dubai crude oil futures, a well-established benchmark for crude oil. Finally, we also consider a strategy involving all the four futures contracts: Brent, Dubai, Shanghai, and WTI.
The results of this analysis are reported in Table 8, where, for the sake of brevity, we fix the bandwidth level $\alpha$ at $0.20$ and omit the $p$-values associated with the [Ledoit and Wolf, 2008] test for the Sharpe ratio and the standard $t$-test for the mean trading-window returns, indicating statistical significance using the conventional notation instead. For comparison purposes, the first row replicates the data from Table 5, while the performances of the benchmark passive strategies, S&P ($R$: 11.74%, $SR$: 0.6729) and ETF ($R$: -28.47%, $SR$: -0.7436), are omitted. Transaction costs specific to each futures contract are again included, as specified in Table 1. Interestingly, when evaluating potential pairs of securities for our strategies, some combinations did not exhibit statistical evidence of cointegration [^9] As explained in Section 4.1.2, cointegration is initially tested using the Johansen methodology. When this test did not detect cointegration, we also applied the Engle-Granger two-step procedure, considering both choices of dependent variable. Even with this alternative method, no evidence of cointegration was found., at least over our training sample (these pairs include WTI-Dubai, Shanghai-WTI and, perhaps unexpectedly, Brent-WTI). This suggests that including at least three contracts in the strategy can be advantageous when standard cointegration does not hold between individual pairs.
As shown in Table 8, the statistical arbitrage strategies based on our modeling framework, namely the PredI and PI strategies, exhibit consistently strong performance across the different sets of securities considered. According to the White Reality Check of [Sullivan et al., 1999], one of these two strategies is always identified as the best-performing alternative. Moreover, the majority of the reported performance measures are statistically significant, with the PredI strategy in particular delivering robust and economically meaningful results across all configurations.
PV ProbI PredI RI PI
@@
$R$ -0.0091 0.1043 0.1518 -0.0244 0.023
SRSR -0.0025 0.8335 1.1792 -0.1508 0.2359
mean $R^{\mathrm{TW}}$ -0.0002 0.0042∗∗ 0.012∗∗ -0.0008 0.0007
$p$-val. WRC 0.065 best from WRC: PredI
@@
$R$ 0.1316 0.104 0.3621 0.0359 0.1711
SRSR 0.7561 0.6382 1.7738∗∗∗ 0.2856 0.9302
mean $R^{\mathrm{TW}}$ 0.0041 0.0044 0.0225∗∗∗ 0.0014 0.0054
$p$-val. WRC 0.011 best from WRC: PredI∗∗
Brent - Dubai
$R$ 0.1478 0.099 0.1214 0.0692 0.1479
SRSR 1.8001∗∗∗ 1.3823∗∗ 1.5267∗∗∗ 1.1975 1.8156∗∗∗
mean $R^{\mathrm{TW}}$ 0.005∗∗∗ 0.0056∗∗∗ 0.007∗∗∗ 0.0032∗∗ 0.0058∗∗∗
$p$-val. WRC 0.0035 best from WRC: PI∗∗∗
Brent - Shanghai - Dubai
$R$ 0.1014 0.0329 0.0692 0.0014 0.1129
SRSR 1.229∗∗ 0.5189 0.915∗∗ 0.053 1.3569∗∗
mean $R^{\mathrm{TW}}$ 0.003∗∗ 0.0019 0.0056∗∗ 0.0001 0.0041∗∗∗
$p$-val. WRC 0.0135 best from WRC: PI∗∗
@@
$R$ 0.113 0.069 0.122 0.0511 0.1163
SRSR 1.4469∗∗ 1.007 1.5607∗∗∗ 0.8929 1.4864∗∗
mean $R^{\mathrm{TW}}$ 0.0045∗∗ 0.0042∗∗ 0.007∗∗∗ 0.0026 0.0053∗∗∗
$p$-val. WRC 0.009 best from WRC: PredI∗∗∗
Brent - Shanghai - WTI - Dubai
$R$ 0.1206 0.0554 0.0636 0.0085 0.1343
SRSR 1.4311∗∗∗ 0.7912 0.8691 0.1657 1.5781∗∗∗
mean $R^{\mathrm{TW}}$ 0.0035∗∗ 0.0032 0.0042 0.0004 0.0045∗∗∗
$p$-val. WRC 0.004 best from WRC: PI∗∗∗
Table 8. Performances over the test sample including different combinations of the Brent, the Dubai, the Shanghai and the WTI futures contracts.